# Multi-task retriever fine-tuning for domain-specific and efficient RAG

### v1

@@
Abstract
Retrieval-Augmented Generation (RAG) has become ubiquitous when deploying Large Language Models (LLMs), as it can address typical limitations such as generating hallucinated or outdated information. However, when building real-world RAG applications, practical issues arise. First, the retrieved information is generally domain-specific. Since it is computationally expensive to fine-tune LLMs, it is more feasible to fine-tune the retriever to improve the quality of the data included in the LLM input. Second, as more applications are deployed in the same real-world system, one cannot afford to deploy separate retrievers. Moreover, these RAG applications normally retrieve different kinds of data. Our solution is to instruction fine-tune a small retriever encoder on a variety of domain-specific tasks to allow us to deploy one encoder that can serve many use cases, thereby achieving low-cost, scalability, and speed. We show how this encoder generalizes to out-of-domain settings as well as to an unseen retrieval task on real-world enterprise use cases.
Multi-task retriever fine-tuning for domain-specific and efficient RAG
Patrice Béchard ServiceNow patrice.bechard@servicenow.com                        Orlando Marquez Ayala ServiceNow orlando.marquez@servicenow.com
1 Introduction
As more and more Generative AI (GenAI) applications are integrated into real-world production systems, Retrieval-Augmented Generation (RAG) has been adopted in industry as a common technique to improve the output of Large Language Models (LLMs). RAG alleviates inherent LLM pitfalls such as propensity to hallucinate, generating outdated knowledge, and lack of traceability to data sources Fan et al. (2024); Gao et al. (2024).
Introducing a retrieval step into the generation process introduces, however, several practical challenges. While an LLM with a large number of parameters, such as GPT-4 OpenAI et al. (2024), can be prompted to work with any kind of input and generate any kind of textual output, the retriever needs to be small, fast, and perform well with data sources that tend to be domain-specific.
Figure 1: Given an ecosystem of RAG applications, how do we build a retriever that can adapt to a specific domain and to a variety of retrieval tasks?
@@ 1 Introduction
Another practical challenge is achieving scalability and generalization across different GenAI use cases that depend on retrieval. A crucial advantage of LLMs compared to traditional machine learning models is that they can generalize to a myriad of tasks due to vast amounts of pretraining data and instruction fine-tuning Wei et al. (2022); Zhang et al. (2024a); Ouyang et al. (2022). But if the retriever does not perform well and fast across many retrieval tasks, the downstream generation will be negatively affected.
@@
The context is an enterprise company that deploys several GenAI applications that currently rely or will rely on RAG: Flow Generation Bechard and Ayala (2024), Playbook Generation222www.servicenow.com/docs/bundle/xanadu-intelligent-experiences/page/administer/now-assist-platform/concept/now-assist-playbook-generation-skill.html, and Code Generation. Workflows are step-by-step processes that automate one or more goals while playbooks contain workflows and other UI components such as forms. The objective of these applications is to generate domain-specific workflows, playbooks, and code from textual input.
@@
To build a multi-task retriever dataset, we extracted data from internal databases and reused the Flow Generation training set. For our solution, we fine-tune mGTE Zhang et al. (2024c) because of its large context length (8,192 tokens), allowing it to receive long instructions, and because of its multilingual capabilities. We compare it with BM25, a simple yet powerful term-frequency method Robertson and Walker (1994), and with recent open-source multilingual embedding models: mE5 Wang et al. (2024) and mGTE.
We perform several evaluations. First, we evaluate on the tasks that the retriever was trained on, but on out-of-domain (OOD) settings. The internal training datasets come from the IT domain, but the OOD splits come from diverse domains such as HR and finance. We then evaluate on a related but different retrieval task to test the generalization ability of the model. For example, while the model was trained to retrieve step names based on text, we also want to retrieve relevant workflow structures based on text. This is a related task because workflows are made up of steps. Lastly, to see whether the multilingual abilities of mGTE are preserved after our multi-task fine-tuning, we evaluate on input from different languages even though our retrieval training dataset is only in English.
@@
2 Related Work
There have been several efforts to achieve domain-specific RAG. While more complex, jointly training the retriever and the generator has been shown to work well for question-answering Siriwardhana et al. (2023). Others encode the domain-specific information in knowledge graphs in addition to vector databases Barron et al. (2024). Retrieval Augmented Fine Tuning Zhang et al. (2024b) helps the LLM adapt to the domain by modifying how the retrieved information is present in the LLM training dataset. In contrast, our approach relies solely on training a better retriever, which can be used with any LLM, thereby reducing coupling.
Multi-task retriever models Maillard et al. (2021); Wang et al. (2022) enhance embedding versatility by simultaneously training on multiple tasks or by utilizing a large dataset spanning diverse topics. Building upon this concept, instruction-following embedding models Asai et al. (2023); Su et al. (2022); BehnamGhader et al. (2024) integrate explicit instructions into the embedding process, thereby enabling the generation of more nuanced and task-specific embeddings. We leverage this kind of instruction fine-tuning to train a retriever on many datasets containing structured data extracted from databases.
Lastly, generating workflows is a structured output task, similar to code generation, that requires specialized embeddings. Code embedding models Feng et al. (2020); Li et al. (2022); Guo et al. (2020); Neelakantan et al. (2022); Zhang* et al. (2024) adapt common training techniques to the domain of code representation learning and retrieval. A parallel research trajectory focuses on structured data embedding models, including Synchromesh Poesia et al. (2022), which fine-tunes a model to retrieve relevant samples for few-shot prompting using a similarity metric derived from tree edit distance, and SANTA Li et al. (2023), which implements a modified pretraining objective to generate structure-aware representations. Prior work already uses RAG for a similar but more limited structured output task Bechard and Ayala (2024). This paper is a natural extension to increase reusability of the retriever in an expanding ecosystem of GenAI applications.
3 Methodology
To build a mult-task retriever, we first had to define the tasks and then build their datasets, keeping in mind that we could not do any labeling for these tasks. The starting point was to define the data that the existing RAG applications would need.
@@ 3 Methodology
Figure 2: Dataset example of Flow Generation. We constructed retrieval multi-task examples from them.
3.1 Tasks
@@ 3.2 Dataset Generation
Following established practices in sentence embedding model training Reimers and Gurevych (2019), we create both positive and negative pairs, enabling the model to effectively discriminate and retrieve relevant elements.
@@ 4.1 Datasets
Train Dev Dev Dev
Pairs Steps Tables Fields
@@
Table 1: Number of training pairs and examples in the development dataset.
@@
4.2 Models
The simplest baseline we compare against is BM25, the simplest in terms of RAG deployment. Stronger baselines are open-source multilingual models. We select the small (118M), base (278M), and large (560M) versions of mE5 Wang et al. (2024) to see whether more parameters help retrieval. Since mE5 has a context length of 512, we also pick mGTE-base Zhang et al. (2024c), which has 8,192 context length and 305M parameters.
Our solution is to fine-tune mGTE-base using a contrastive loss objective Hadsell et al. (2006). We prefer a retriever with large context as, once deployed, we do not control the length of instructions/context that it will receive. We also prefer a small model to allow the retrieval part of RAG to scale to many LLM concurrent calls. Appendix A provides details of our training process, including hyperparameters.
4.3 Evaluation Metrics
We use Recall@K Manning (2008) as our metric across all tasks. While we briefly explored more sophisticated metrics such as Normalized Discounted Cumulative Gain (NDCG) Järvelin and Kekäläinen (2002) and Mean Reciprocal Rank (MRR) Voorhees et al. (1999), we found that these metrics highly correlate with recall in our use case.
@@ 5 Results
5.1 Multi-task Retriever
@@
Table 3: Single task vs Multi-task and effect of balanced dataset. Evaluation on development split.
@@
5.2 Comparison to Baselines on OOD Splits
@@
Ours 0.90 0.90 0.60
Table 4: Performance of BM25, open source models, and multi-task instruction fine-tuned mGTE-base (ours). Results are weighted averages across all OOD splits.
On our domain-specific retrieval tasks, our multi-task fine-tuned model surpasses BM25, and the mE5 and mGTE open source alternatives. It is notable that BM25 beats all the considered open source models. Scaling the base retriever from mE5-small to mE5-large did not improve results.
@@
5.3 Effect on Multilingual Capabilities
@@
Table 5: Results on multilingual development datasets across step, table, and field retrieval.
@@
5.4 Generalization to Workflow Retrieval
@@
Table 6: Generalization on workflow retrieval task. Metric is recall@5.
6 Conclusion
@@ Appendix A Training Details
We fine-tuned mGTE-base for 5,000 steps, equivalent to approximately 2 epochs on the dataset. Training hyperparameters were as follows: a batch size of 32 was employed, with a per-device batch size of 2 and 16 gradient accumulation steps. Since we are not using in-batch negatives, we do not need to employ techniques such as GradCache Gao et al. (2021) to account for larger batch sizes. The learning rate was set to 5e-5, with a weight decay of 0.01. A warmup period of 500 steps was implemented, followed by a cosine learning rate scheduler. To optimize memory usage, gradient checkpointing was utilized Chen et al. (2016). The Adafactor optimizer Shazeer and Stern (2018) was chosen for limited memory consumption.

### v2

@@
Abstract.
Retrieval-Augmented Generation (RAG) has become ubiquitous when deploying Large Language Models (LLMs), as it can address typical limitations such as generating hallucinated or outdated information. However, when building real-world RAG applications, practical issues arise. First, the retrieved information is generally domain-specific. Since it is computationally expensive to fine-tune LLMs, it is more feasible to fine-tune the retriever model to improve the quality of the data included in the LLM input. Second, as more applications are deployed in the same real-world system, one cannot afford to deploy separate retrievers for different tasks. Moreover, these RAG applications normally retrieve different kinds of data. Our solution is to instruction fine-tune a small retriever model on a variety of domain-specific tasks to allow us to deploy one model that can serve many use cases, thereby achieving low-cost, scalability, and speed. We show how this model generalizes to out-of-domain settings as well as to an unseen retrieval task on real-world enterprise use cases.
Retrieval-Augmented Generation, Workflows, Fine-Tuning, Multi-Task, Retriever
copyright: noneconference: Workshop on Structured Knowledge for Large Language Models (SKnowLLM) at KDD 2025; August 03–07,2025; Toronto, Canadaccs: Computing methodologies Information extractionccs: Computing methodologies Multi-task learningccs: Computing methodologies Learning to rankccs: Information systems Retrieval models and ranking
1. Introduction
As more and more Generative AI (GenAI) applications are integrated into real-world production systems, Retrieval-Augmented Generation (RAG) has been adopted in industry as a common technique to improve the output of Large Language Models (LLMs). RAG alleviates inherent LLM pitfalls such as propensity to hallucinate, generating outdated knowledge, and lack of traceability to data sources (Fan et al., 2024; Gao et al., 2024).
Introducing a retrieval step into the generation process introduces, however, several practical challenges. While an LLM with a large number of parameters, such as GPT-4 (OpenAI et al., 2024), can be prompted to work with any kind of input and generate any kind of textual output, the retriever needs to be small, fast, and perform well with data sources that tend to be domain-specific.
Figure 1. Given an ecosystem of RAG applications, how do we build a retriever that can adapt to a specific domain and to a variety of retrieval tasks?
[…]
Architecture of structured data retrieval for multiple tasks.
@@ 1. Introduction
Another practical challenge is achieving scalability and generalization across different GenAI use cases that depend on retrieval. A crucial advantage of LLMs compared to traditional machine learning models is that they can generalize to a myriad of tasks due to vast amounts of pretraining data and instruction fine-tuning (Wei et al., 2022; Zhang et al., 2024a; Ouyang et al., 2022). But if the retriever does not perform well and fast across many retrieval tasks, the downstream generation will be negatively affected.
@@
The context is an enterprise company that deploys several GenAI applications that currently rely or will rely on RAG: Flow Generation (Bechard and Ayala, 2024), Playbook Generation222www.servicenow.com/docs/bundle/xanadu-intelligent-experiences/page/administer/now-assist-platform/concept/now-assist-playbook-generation-skill.html, and Code Generation. Workflows are step-by-step processes that automate one or more goals while playbooks contain workflows and other UI components such as forms. The objective of these applications is to generate domain-specific workflows, playbooks, and code from textual input.
@@
To build a multi-task retriever dataset, we extracted data from internal databases and reused the Flow Generation training set. For our solution, we fine-tune mGTE (Zhang et al., 2024c) because of its large context length (8,192 tokens), allowing it to receive long instructions, and because of its multilingual capabilities. We compare it with BM25, a simple yet powerful term-frequency method (Robertson and Walker, 1994), and with recent open-source multilingual embedding models: mE5 (Wang et al., 2024b) and mGTE. We also include evaluation of larger instruction-tuned models based on LLMs, namely GTE-Qwen-2 (Yang et al., 2024), E5-mistral (Wang et al., 2024a), and GritLM (Muennighoff et al., 2024).
We perform several evaluations. First, we evaluate on the tasks that the retriever was trained on, but on out-of-domain (OOD) settings. The internal training datasets come from the IT domain, but the OOD splits come from diverse domains such as HR and finance. We then evaluate on a related but different retrieval task to test the generalization ability of the model. For example, while the model was trained to retrieve step names based on text, we also want to retrieve relevant workflow structures based on text. This is a related task because workflows are made up of steps. Lastly, to see whether the multilingual abilities of models are preserved after our multi-task fine-tuning, we evaluate on input from different languages even though our retrieval training dataset is only in English.
@@
2. Related Work
There have been several efforts to achieve domain-specific RAG. While more complex, jointly training the retriever and the generator has been shown to work well for question-answering (Siriwardhana et al., 2023). Others encode the domain-specific information in knowledge graphs in addition to vector databases (Barron et al., 2024). Retrieval-Augmented Fine Tuning (Zhang et al., 2024b) helps the LLM adapt to the domain by modifying how the retrieved information is present in the LLM training dataset. In contrast, our approach relies solely on training a better retriever, which can be used with any LLM, thereby reducing coupling.
Multi-task retriever models (Maillard et al., 2021; Wang et al., 2022) enhance embedding versatility by simultaneously training on multiple tasks or by utilizing a large dataset spanning diverse topics. Building upon this concept, instruction-following embedding models (Asai et al., 2023; Su et al., 2022; BehnamGhader et al., 2024) integrate explicit instructions into the embedding process, thereby enabling the generation of more nuanced and task-specific embeddings. We leverage this kind of instruction fine-tuning to train a retriever on many datasets containing structured data extracted from databases.
Lastly, generating workflows is a structured output task, similar to code generation, that requires specialized embeddings. Code embedding models (Feng et al., 2020; Li et al., 2022; Guo et al., 2020; Neelakantan et al., 2022; Zhang* et al., 2024) adapt common training techniques to the domain of code representation learning and retrieval. A parallel research trajectory focuses on structured data embedding models, including Synchromesh (Poesia et al., 2022), which fine-tunes a model to retrieve relevant samples for few-shot prompting using a similarity metric derived from tree edit distance, and SANTA (Li et al., 2023), which implements a modified pretraining objective to generate structure-aware representations. Prior work already uses RAG for a similar but more limited structured output task (Bechard and Ayala, 2024). This paper is a natural extension to increase reusability of the retriever in an expanding ecosystem of GenAI applications.
3. Methodology
To build a multi-task retriever, we first had to define the tasks and then build their datasets, keeping in mind that we could not do any labeling for these tasks. The starting point was to define the data that the existing RAG applications would need.
@@ 3. Methodology
Figure 2. Dataset example of Flow Generation. We constructed retrieval multi-task examples from them.
[…]
Example of workflow JSON
3.1. Tasks
@@ 3.2. Dataset Generation
Following established practices in sentence embedding model training (Reimers and Gurevych, 2019), we create both positive and negative pairs, enabling the model to effectively discriminate and retrieve relevant elements.
@@ 4.1. Datasets
Table 1. Number of training pairs and examples in the development dataset.
Train Pairs Dev Steps Dev Tables Dev Fields
@@
4.2. Models
The simplest baseline we compare against is BM25, the simplest in terms of RAG deployment. Stronger baselines are open-source multilingual models. We select the small (118M), base (278M), and large (560M) versions of mE5 (Wang et al., 2024b) to see whether more parameters help retrieval. Since mE5 has a context length of 512, we also pick mGTE-base (Zhang et al., 2024c), which has 8,192 context length and 305M parameters.
Our solution is to fine-tune mGTE-base using a contrastive loss objective (Hadsell et al., 2006). We prefer a retriever with large context as, once deployed, we do not control the length of instructions/context that it will receive. We also prefer a small model to allow the retrieval part of RAG to scale to many LLM concurrent calls. Appendix A provides details of our training process, including hyperparameters.
4.3. Evaluation Metrics
We use Recall@K (Manning, 2008) as our metric across all tasks. While we briefly explored more sophisticated metrics such as Normalized Discounted Cumulative Gain (NDCG) (Järvelin and Kekäläinen, 2002) and Mean Reciprocal Rank (MRR) (Voorhees et al., 1999), we found that these metrics highly correlate with recall in our use case.
@@ 5. Results
5.1. Multi-task Retriever
@@
Table 3. Single task vs Multi-task and effect of balanced dataset. Evaluation on development split.
@@
5.2. Comparison to Baselines on OOD Splits
@@
Table 4. Performance of BM25, small and large open source models, and multi-task instruction fine-tuned mGTE-base (ours). Results are weighted averages across all OOD splits.
@@
GTE-Qwen2-1.5B-Instruct 0.21 0.12 0.01
GTE-Qwen2-7B-Instruct 0.18 0.08 0.03
GritLM-7B 0.75 0.59 0.16
E5-Mistral-7B-Instruct 0.57 0.51 0.16
Finetuned mGTE-Base (Ours) 0.90 0.90 0.60
On our domain-specific retrieval benchmarks, our multi-task fine-tuned mGTE-base model achieves the highest performance across all metrics, substantially outperforming both BM25 and all evaluated open-source embedding models. Notably, BM25 consistently outperforms the open-source embedding models — regardless of model size — including larger variants such as mE5-large and GTE-Qwen2. Increasing the scale of the mE5 retrievers from small to large yields no improvement, and in some cases results in lower scores. Only our model delivers strong, consistent gains, reaching 0.90 on both Step@15 and Table@5, and 0.60 on Field@5.
@@
5.3. Effect on Multilingual Capabilities
@@
Table 5. Results on multilingual development datasets across step, table, and field retrieval.
@@
5.4. Generalization to Workflow Retrieval
@@
Table 6. Generalization on workflow retrieval task. Metric is recall@5.
@@
6. Conclusion
@@ Appendix A Training Details
We fine-tuned mGTE-base for 5,000 steps, equivalent to approximately 2 epochs on the dataset. Training hyperparameters were as follows: a batch size of 32 was employed, with a per-device batch size of 2 and 16 gradient accumulation steps. Since we are not using in-batch negatives, we do not need to employ techniques such as GradCache (Gao et al., 2021) to account for larger batch sizes. The learning rate was set to 5e-5, with a weight decay of 0.01. A warmup period of 500 steps was implemented, followed by a cosine learning rate scheduler. To optimize memory usage, gradient checkpointing was utilized (Chen et al., 2016). The Adafactor optimizer (Shazeer and Stern, 2018) was chosen for limited memory consumption.